Consider being dealt 7 cards from a standard deck of 52. What is the probability that you will have exactly 4 spades?
Answer by venugopalramana(3286)

Consider being dealt 7 cards from a standard deck of 52. What is the probability that you will have exactly 4 spades?
THERE ARE 13 SPADES AND 39 OTHER CARDS IN A PACK.
4 SPADES CAN BE GOT FROM 13 SPADES IN 13C4 WAYS.=13!/4!*(13-4)!=13*12*11*10/(4*3*2*1)=13*11*5
3 OTHER CARDS GOT FROM 39 OTHER CARDS IN 39C3 WAYS=39!/3!*(39-3)!=39*38*37/(3*2*1)=13*19*37
TOTAL NO. OF 7 CARD COMBINATIONS ARE = 13*11*5*13*19*37